lim (sin2x) tan^2(2x) x_pie/4

lim        (sin2x)tan^2(2x)
x_pie/4

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
1infiniteform
lim x_pie/4 (1+sin2x-1)^tan^2(2x)=e^(lim x_pie/4 (sin2x-1)*(tan^2(2x)))
apply L Hospital rule because 0/0 form
e^(lim x_pie/4 (sin2x-1)/(cot^2(2x))
=e^(lim x_pie/4 (2*cos2x)/(4*cot2x*cosec^2(2x))
=e^(lim x_pie/4 (2*cos2x*sin^3(2x))/(4*cos2x)
=e^(1/2)

Sher Mohammad
faculty askiitians